C=100+0.8yd G=100 T=0.25y X=150 M=0.25yd I=200 1) What is the level of equilibrium national income? 2) Estimate the budget surplus or deficit at the equilibrium national income. 3) Derive the savings function. Interpret your answer. 4) Define the income multiplier (=k).
1
Expert's answer
2012-08-09T08:31:41-0400
1) Yd= Y - T
AE = C + I + G + (X - M)
AE = Y
Y=100+0,8(Y-T)+100+200+150-0,25(Y-T)
Y=100+0,6Y+100+200+150-0,1875Y
Ye=936,2 (the level of equilibrium national income)

2) Budget surplus is when X>M, otherwise (X<M) there is a budget deficit. So at the equilibrium national income there is a
deficit (150<175,5).

3) S = Y - C , where Saving = Income - Consumption
S = Y - (100 + 0,8Yd)
S = 1397,9

4) k = 1/ (1 - MPC)
k = 16,7